# Materialist: Physically Based Editing Using Single-Image Inverse Rendering

### v1

@@
Abstract
To perform image editing based on single-view, inverse physically based rendering, we present a method combining a learning-based approach with progressive differentiable rendering. Given an image, our method leverages neural networks to predict initial material properties. Progressive differentiable rendering is then used to optimize the environment map and refine the material properties with the goal of closely matching the rendered result to the input image. We require only a single image while other inverse rendering methods based on the rendering equation require multiple views. In comparison to single-view methods that rely on neural renderers, our approach achieves more realistic light material interactions, accurate shadows, and global illumination. Furthermore, with optimized material properties and illumination, our method enables a variety of tasks, including physically based material editing, object insertion, and relighting. We also propose a method for material transparency editing that operates effectively without requiring full scene geometry. Compared with methods based on Stable Diffusion, our approach offers stronger interpretability and more realistic light refraction based on empirical results. Code is avaiable at github.com/lez-s/Materialist
@@ Abstract
Figure 1: Given an image (in the middle), our inverse rendering approach enables physically based image editing, including transparency, albedo, roughness, metallic, transparent object insertion and relighting. The image is sourced from the IIW dataset [4].
1 Introduction
@@
High-quality image editing often requires professional skills. Reducing the complexity and increasing the accuracy of image editing has long been a focus in computer vision and computer graphics [36, 40, 45, 3]. With the success of Stable Diffusion (SD) [48] in image generation, researchers have explored using SD for image editing [19, 38, 7, 77]. However, SD often struggles with precise material property editing. Alchemist [53] attempts to address this by training SD on a large synthetic dataset, but their method lacks interpretability, a physical foundation, and struggles with accurate light refraction for transparent objects.
@@
Inverse rendering offers a promising solution to these challenges. Single-view inverse rendering [28, 29, 79, 52] decomposes images into albedo, roughness, and metallic properties via neural networks (MaterialNet), enhancing interpretability. However, relying on a neural renderer as a substitute for a physically based renderer still shares limitations with SD, such as difficulties in accurately handling of shadows and refraction. Multi-view inverse rendering [67, 1, 69, 71, 30, 2] leverages multi-view constraints and the rendering equation to adhere better to physical principles and achieve superior results. The multi-view input requirement, however, restricts its applicability. Applying differentiable rendering to a single image leads to an ill-posed problem with infinitely many possible solutions.
@@
To address these issues, we adopt a more interpretable inverse rendering approach that combines the strengths of single- and multi-view methods (MaterialNet and the rendering equation). Additionally, we introduce material transparency editing, a capability not previously explored in single-image inverse rendering.
Our main contributions are threefold:
@@
  • Physically Based Material Editing: Instead of a neural renderer, our method leverages Mitsuba [20] to achieve physically based material editing with precise global illumination, shadows, and realistic object-environment interactions, thus outperforming existing methods in material editing, relighting, and transparent object insertion tasks.
  • Single-View Transparency Editing: We propose a single-view, physically based transparency editing method with index of refraction (IOR) and specular transmission, which operates under the single-view constraint without requiring complete geometry or full scene reconstruction.
In addition, we demonstrate that our material prediction model outperforms the latest model based on Stable Diffusion [24] with much faster inference. We also achieve better results in single-view differentiable Monte Carlo ray tracing for environment maps and introduce a mesh reconstruction method well-suited for single-image inverse rendering.
2 Related work
Figure 2: Our inverse rendering pipeline. Given an image, we use MatNet to predict material properties, followed by envmap optimization to estimate the lighting. We then perform material properties optimization, using the envmap that yields the smallest $\mathcal{L}_{\text{re}}$ during envmap optimization as the light source. The losses $\mathcal{L}_{\text{re}}$ and $\mathcal{L}_{\text{cons}}$ guide this process, allowing $\mathcal{L}_{\text{re}}$ to be minimized while keeping the results close to the MatNet predictions. See Sec. 3.2 for further details.
@@ 2 Related work
Early work like IPT [1] was limited to simple scenes. They used multi-view constraints and assumed known scene geometry when using the rendering equation for inverse rendering. With NeRF [37], physically based inverse rendering with multi-view images became more feasible [69, 71, 30, 2, 56]. In addition, use of differentiable Monte Carlo rendering led to better shape and material estimation [33, 18, 76, 67, 73, 64], but these techniques still rely on multi-view inputs. More recent methods employ neural networks as renderers for multi-view inputs [12], bypassing computational optimization but currently lacking the performance of equation-based methods in material editing and object insertion. Like most multi-view techniques, our approach uses the rendering equation but requires only a single-view image as input.
@@
Unlike multi-view inverse rendering, which relies on the rendering equation, single-view inverse rendering often uses neural networks trained on large datasets to replace physically based rendering [26, 27, 29, 49, 52, 79, 34]. A neural network usually called MaterialNet, predicts material properties (albedo, roughness, metallic) and per-pixel lighting from a single image, which are then used for image synthesis. Although neural renderers perform well in specific tasks like human face relighting [70, 41], they struggle with generalist tasks where physical simulation is more accurate such as material editing and transparent object insertion. Unless trained on a well represented dataset, the neural renderers fall apart.
In differentiable rendering, early studies addressed rasterization non-differentiability [11, 10], while recent work has explored inverse Monte Carlo ray tracing [21, 75, 14], similar to our physics-based optimization. However, these methods require significant priors and are limited to single objects, making them less suitable for complex scenes. Our approach also uses MatNet for single-view material property inference but incorporates a physics-based renderer.
@@
Material editing generally follows two categories. The first involves inferring material properties through inverse rendering, then modifying them to achieve edits [25, 78, 1]. These methods provide strong interpretability, and using a physics-based renderer makes results highly reliable, though realism may sometimes be limited. The second set is neural network-based. Following the success of Stable Diffusion [48], many image editing methods have emerged [38, 63, 7, 55]. Alchemist [53] is a representative example, trained on synthetic data to allow editing of albedo, roughness, metallic properties, and transparency. These methods benefit from the strengths of Stable Diffusion and achieve highly realistic results but have limited interpretability due to their neural network-based nature. Our approach falls into the first category with a solid interpretability, meanwhile, it enables realistic results.
@@
Light estimation is a distinct research field [62, 60, 42, 57, 16, 72, 58, 13]. Some methods represent lighting implicitly [69, 78, 24], limiting generalizability, while others require multi-view inputs and scene mesh data [60, 72, 42]. DPI [35] is related to our approach. This method combines differential rendering with Stable Diffusion for high-quality envmap generation but relies on multi-view NeRF methods for mesh reconstruction and lacks material BRDF data, reducing accuracy. In contrast, our approach uses MatNet’s output for more accurate envmap optimization.
@@
Shadows are crucial for photorealistic rendering. Previous work by Li et al. [29] emphasized shadow accuracy, using an OptiX-based ray tracer [43] to compute shadows for neural network input. However, shadows are often neglected in recent research [78, 69, 30]; methods based on screen-space or image-based ray tracing lack geometric occlusion and thus fail to generate accurate shadows [69, 24, 78]. Many neural renderers struggle with shadow generation [12, 79], as accurate shadows require physically based light transport, it is worth noting that this limitation has been utilized to identify AI-generated images [50]. Our method includes full path tracing for physically accurate shadows.
3 Method
Our material estimation and editing framework consists of four main steps. For relighting, only the geometry reconstruction and the material prediction network (Sec. 3.1) are needed. Object insertion and material editing on the other hand require all four steps, including mesh reconstruction. We assume a pinhole camera model with a 35-degree field of view, $w\times{h}=512\times 512$ image resolution, and $c=3$ color channels.
3.1 Material Prediction Network
@@
Due to the significant success of Stable Diffusion [48] in image generation, recent work used it for the task of material property estimation, achieving impressive results [24]. However, this approach comes with high training costs and requires computationally intensive multi-step sampling during prediction. In comparison, we found that the model architecture for depth estimation [68] aligns more closely with the requirements of material prediction. We therefore adopt the DPT architecture [46], which has been highly successful in depth estimation, for our material prediction task. Different from [79], we utilize a pretrained DINOv2 [39] encoder for feature extraction. Like previous research in depth estimation [6, 68], we use two DPT decoders [46] for the regression of depth and material properties, respectively. Given the advancements in depth estimation, we initialize our training with weights from the pretrained depth model [68] to expedite the training process. For the DPT depth decoder, we limit training to the last four layers of the refinenet, while the weights of the DINOv2 encoder remain frozen throughout the training. The following loss function for $\mathcal{L}_{\text{MatNet}}$ is used to optimize the model,
$\mathcal{L}_{\text{MatNet}}=\sum_{i\in\{\mathbf{A},\mathbf{R},\mathbf{M},\mathbf{N},\mathbf{D}\}}\mathcal{L}_{i}\,,$ (1)
@@
where the albedo loss $\mathcal{L}_{\mathbf{A}}$ combines the LPIPS perceptual loss and $L_{1}$ loss, while both the roughness loss $\mathcal{L}_{\mathbf{R}}$ and metallic loss $\mathcal{L}_{\mathbf{M}}$ are measured using the $L_{1}$ loss between the predictions and ground truth values. Finally, the $\mathcal{L}_{\mathbf{D}}$ is calculated using SiLogLoss [5, 15] to introduce logarithmic error and keep its scale constant to improve the robustness of depth prediction. This loss is defined by
@@
$\mathcal{L}_{\mathbf{D}}=\sqrt{\begin{aligned} &\mathbb{E}_{i\in\text{mask}}\left(\left(\log(t_{i})-\log(p_{i})\right)^{2}\right)\\ &-\lambda\left(\mathbb{E}_{i\in\text{mask}}\left(\log(t_{i})-\log(p_{i})\right)\right)^{2}\,,\end{aligned}}$ (2)
@@
$\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p},\mathbf{N}_{p},\mathbf{D}_{p}=\text{MatNet}(I)\,,$ (3)
@@
Figure 3: Our single-view mesh reconstruction method used on a test scene. Our method eliminates the white border at the edges. However, since our approach is based on the depth map, it primarily reconstructs the front view of the scene.
@@
3.2 Progressive Differentiable Rendering
Our method assumes ambient lighting, represented by an environment map, and does not explicitly model light sources within the scene. Ignoring emitters in the scene, the rendering equation becomes the reflected radiance equation:
@@ 3.2 Progressive Differentiable Rendering
where $L_{o}(\bm{x},\vec{\omega}_{o})$ is the outgoing radiance at the world space position $\bm{x}$ in the direction $\vec{\omega}_{o}$, $f_{r}(\bm{x},\vec{\omega}_{i},\vec{\omega}_{o})$ is the bidirectional reflectance distribution function (BRDF) at $\bm{x}$, representing the ratio of reflected radiance in the direction $\vec{\omega}_{o}$ to irradiance from the direction $\vec{\omega}_{i}$, while $L_{i}(\bm{x},\vec{\omega}_{i})$ is the incident radiance at $\bm{x}$ from the direction $\vec{\omega}_{i}$, $\vec{n}$ is the surface normal at $\bm{x}$, and $\Omega$ is the hemisphere around $\vec{n}$ covering all directions of incidence. We let $\mathbf{S}(\bm{x})=\bm{x}_{s}$ denote the transformation from world space to screen space coordinates. Detailed steps are in the supplementary material.
Simplified DisneyBRDF.
To model surface scattering we use the DisneyBRDF [8, 9] for our implementation. During the optimization process, we do not model sheen, clearcoat, and glass, so the BRDF in Eq. 4 can be written
$f_{r}(\bm{x},\vec{\omega}_{i},\vec{\omega}_{o})=f_{s}+(1-\mathbf{M}(\bm{x}_{s}))f_{d}$ (5)
with the specular ($f_{s}$) and diffuse ($f_{d}$) terms defined by
$f_{s}$ $=\frac{F_{s}(\vec{\omega}_{i},\vec{\omega}_{o};\eta)D_{s}(\vec{h};\mathbf{R}(\bm{x}_{s}))G_{s}(\vec{\omega}_{i},\vec{\omega_{o}};\mathbf{R}(\bm{x}_{s}))}{4\,|\vec{n}\cdot\vec{\omega}_{i}||\vec{n}\cdot\vec{\omega}_{o}|}$ (6)
$f_{d}$ $=\frac{\mathbf{A}(\bm{x}_{s})}{\pi}F_{d}(\vec{\omega}_{i})F_{d}(\vec{\omega}_{o})\,,$ (7)
where $F$ is the Fresnel reflectance, which here uses the half vector $\vec{h}=(\vec{\omega}_{i}+\vec{\omega}_{o})/\|\vec{\omega}_{i}+\vec{\omega}_{o}\|$ and models the fraction of light reflected from the surface as a function of material properties, including the relative index of refraction $\eta$. We use the modified Schlick approximation [8]:
$\!\!F_{s}(\vec{\omega}_{i},\vec{\omega}_{o};\eta)$ $=C_{0}(\eta)+(1-C_{0}(\eta))(1-\vec{h}\cdot\vec{\omega}_{o})^{5}\,$ (8)
$C_{0}(\eta)$ $=R_{0}(\eta)(1-\mathbf{M}(\bm{x}_{s}))+\mathbf{M}(\bm{x}_{s})\mathbf{A}(\bm{x}_{s})\,,$ (9)
$F_{d}(\vec{\omega})$ $=1+(F_{D90}-1)(1-\vec{n}\cdot\vec{\omega})^{5}\,$ (10)
$F_{D90}$ $=0.5+2\,\mathbf{R}(\bm{x}_{s})\,(\vec{h}\cdot\vec{\omega})^{2}\,.$ (11)
Regarding the other terms, $D_{s}$ is the microfacet normal distribution function, while $G_{s}$ is the microfacet shadowing-masking function [51]. We use the GGX distribution [61] for these terms due to its simplicity, with just one roughness parameter $\mathbf{R}(\bm{x}_{s})$, while being able to fit empirical data well.
@@
We optimize the environment map $\mathbf{E}$ (envmap) using a position-embedded multi-layer perceptron (MLP), similar to the approach used in NeRF [37]. Although optimizing $\mathbf{E}$ directly without using a neural network is theoretically possible, this often yields suboptimal results. Standard MLPs are generally insensitive to high-frequency information, making pixel-level optimization of the image difficult; CNNs encounter similar limitations. The envmap is obtained using
$\mathbf{E}=\mathrm{MLP}_{\text{env}}(\gamma(\mathbf{e});\theta_{\text{env}})\,,$ (12)
@@
In a differentiable rendering of the geometry $\mathbf{D}_{p},\mathbf{N}_{p}$ based on Monte Carlo integration of Eq. 4, we evaluate the BRDF using $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}$ (Eq. 5) and the incident radiance using
$L_{i}(\bm{x},\vec{\omega}_{i})=V(\bm{x},\vec{\omega}_{i})L_{\mathbf{E}}(\vec{\omega}_{i})\,,$ (13)
where $L_{\mathbf{E}}(\vec{\omega}_{i})$ is the radiance from the envmap in the direction $\vec{\omega}_{i}$ and $V(\bm{x},\vec{\omega}_{i})$ is visibility, which is zero if geometry is found in the direction $\vec{\omega}_{i}$ and one otherwise.
Our optimization goal is to find the $\mathbf{E}^{\ast}$ that minimizes the difference between the rendering and the ground truth ($I$):
$\mathbf{E}^{\ast}=\operatorname*{\arg\!\min}_{\mathbf{E}}\mathcal{L_{\mathbf{E}}}(L_{o}(\bm{x},\vec{\omega}_{o}),I)$ (14)
where environment map loss $\mathcal{L_{\mathbf{E}}}$ combines $L_{1}$ and $L_{2}$ loss. The network weights $\theta_{\text{env}}$ are discarded after optimization as we only need $\mathbf{E}^{\ast}$.
@@
$\left[\mathbf{A},\mathbf{R},\mathbf{M}\right]=\mathrm{MLP}_{\text{mat}}(\gamma(\mathbf{m});\theta_{\text{mat}})$ (15)
$\mathbf{m}=\left[\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}\right]\,,$ (16)
@@
$\left[\mathbf{A},\mathbf{R},\mathbf{M}\right]=\zeta\,\tanh(\mathbf{z})+\mathbf{m}\,,$ (17)
@@
The rendering process is the same as for the envmap, except that now we use $\bm{E}^{\ast}$ and keep it fixed.
Our optimization goal is to find $\mathbf{A}^{\ast},\mathbf{R}^{\ast},\mathbf{M}^{\ast}$ that minimize the difference between the rendered result and the ground truth ($I$):
$\mathbf{A}^{\ast},\mathbf{R}^{\ast},\mathbf{M}^{\ast}=\operatorname*{\arg\!\min}_{\mathbf{A,R,M}}\mathcal{L}_{\text{Mat}}(L_{o}(\bm{x},\vec{\omega}_{o}),I)$ (18)
@@
$\mathcal{L}_{\text{Mat}}$ $=\mathcal{L}_{\text{re}}+\delta\mathcal{L}_{\text{cons}}\,,$ (19)

### v2

@@
Abstract.
Achieving physically consistent image editing remains a significant challenge in computer vision. Existing image editing methods typically rely on neural networks, which struggle to accurately handle shadows and refractions. Conversely, physics-based inverse rendering often requires multi-view optimization, limiting its practicality in single-image scenarios. In this paper, we propose Materialist, a method combining a learning-based approach with physically based progressive differentiable rendering. Given an image, our method leverages neural networks to predict initial material properties. Progressive differentiable rendering is then used to optimize the environment map and refine the material properties with the goal of closely matching the rendered result to the input image. Our approach enables a range of applications, including material editing, object insertion, and relighting, while also introducing an effective method for editing material transparency without requiring full scene geometry. Furthermore, Our envmap estimation method also achieves state-of-the-art performance, further enhancing the accuracy of image editing task. Experiments demonstrate strong performance across synthetic and real-world datasets, excelling even on challenging out-of-domain images.
Relight, Physically based rendering, Inverse Rendering, Light estimation, Material estimation
submissionid: 1742copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologiesccs: Computing methodologies Appearance and texture representations
@@ Abstract.
Figure 1. Given an image (in the middle), our inverse rendering approach enables physically based image editing, including transparency, albedo, roughness, metallic, transparent object insertion and relighting. The image is sourced from the IIW dataset (Bell et al., 2014).
[…]
1. Introduction
@@
High-quality image editing always requires professional skills. Reducing the complexity and increasing the accuracy of image editing has long been a focus in computer vision and computer graphics (Meng et al., 2021; Pan et al., 2023; Popat and Picard, 1993; Barnes et al., 2009). With the success of Generative Diffusion Model (DM) (Rombach et al., 2022) in image generation, researchers have explored using DM for image editing (Hertz et al., 2023; Mokady et al., 2023; Brooks et al., 2023). However, DM often struggles with precise material property editing. Alchemist (Sharma et al., 2024) attempts to address this by training DM on a large synthetic dataset, but their method lacks interpretability, a physical foundation, and struggles with accurate light refraction for transparent objects.
@@
Inverse rendering offers a promising solution to these challenges. Single-view inverse rendering (Li et al., 2021, 2022; Zhu et al., 2022a; Sengupta et al., 2019) decomposes images into albedo, roughness, and metallic properties via neural networks (MaterialNet), enhancing interpretability. However, relying on a neural renderer as a substitute for a physically based renderer still shares limitations with DM, such as difficulties in accurately handling of shadows and refraction. Multi-view inverse rendering (Wu et al., 2023; Azinovic et al., 2019; Yao et al., 2022; Yu et al., 2023b; Li et al., 2023; Bao et al., 2024) leverages multi-view constraints and the rendering equation to adhere better to physical principles and achieve superior results. The multi-view input requirement, however, restricts its applicability. Applying differentiable rendering to a single image leads to an ill-posed problem with infinitely many possible solutions.
@@
To address these issues, we adopt a more interpretable inverse rendering approach that combines the strengths of learning and physically based rendering. Additionally, we introduce material transparency editing, a capability not previously explored in single-image inverse rendering.
Our main contributions are:
@@
  • Better Envmap Estimation: Benefiting from MaterialNet’s predictions of material properties, we achieve better results in single-view differentiable Monte Carlo ray tracing for environment maps.
  • Physically Based Material Editing: Instead of a neural renderer, our method leverages Mitsuba (Jakob et al., 2022) to achieve physically based material editing with precise global illumination, shadows, and realistic object-environment interactions, thus outperforming existing methods in material editing, relighting, and transparent object insertion tasks.
  • Single-Image Transparency Editing: We propose a single-view, physically based transparency editing method with index of refraction (IOR) and specular transmission, which operates under the single-view constraint without requiring complete geometry or full scene reconstruction.
In addition, our material prediction model outperforms the latest DM-based model (Kocsis et al., 2024; Zeng et al., 2024) on synthetic datasets, approaches its best performance on real-world IIW datasets (Bell et al., 2014), and runs over 30× faster. We also introduce a mesh reconstruction method well-suited for single-image inverse rendering.
2. Related work
Figure 2. Our inverse rendering pipeline. Given an image, we use MatNet to predict material properties, followed by envmap optimization to estimate the lighting. We then perform material properties optimization, using the envmap that yields the smallest $\mathcal{L}_{\text{re}}$ during envmap optimization as the light source. The losses $\mathcal{L}_{\text{re}}$ and $\mathcal{L}_{\text{cons}}$ guide this process, allowing $\mathcal{L}_{\text{re}}$ to be minimized while keeping the results close to the MatNet predictions. See Sec. 3.2 for further details.
@@ 2. Related work
Early work like IPT (Azinovic et al., 2019) was limited to simple scenes. They used multi-view constraints and assumed known scene geometry when using the rendering equation for inverse rendering. With NeRF (Mildenhall et al., 2021), physically based inverse rendering with multi-view images became more feasible (Yao et al., 2022; Yu et al., 2023b; Li et al., 2023; Bao et al., 2024; Sun et al., 2023). In addition, use of differentiable Monte Carlo rendering led to better shape and material estimation (Luan et al., 2021; Hasselgren et al., 2022; Zhang et al., 2022; Wu et al., 2023; Yu et al., 2022; Wang et al., 2021a), but these techniques still rely on multi-view inputs. More recent methods employ neural networks as renderers for multi-view inputs (Choi et al., 2023), bypassing computational optimization but currently lacking the performance of equation-based methods in material editing and object insertion. Like most multi-view techniques, our approach uses the rendering equation but requires only a single-view image as input.
@@
Unlike multi-view inverse rendering, which relies on the rendering equation, single-view inverse rendering often uses neural networks trained on large datasets to replace physically based rendering (Li et al., 2018b, 2020, 2022; Sang and Chandraker, 2020; Sengupta et al., 2019; Zhu et al., 2022a; Luo et al., 2024). A neural network usually called MaterialNet, predicts material properties (albedo, roughness, metallic) and per-pixel lighting from a single image, which are then used for image synthesis. Although neural renderers perform well in specific tasks like human face relighting (Yeh et al., 2022; Pandey et al., 2021), they struggle with generalist tasks where physical simulation is more accurate such as material editing and transparent object insertion. Unless trained on a well represented dataset, the neural renderers fall apart.
In differentiable rendering, early studies addressed rasterization non-differentiability (Chen et al., 2021, 2019), while recent work has explored inverse Monte Carlo ray tracing (Kato et al., 2020; Zhang et al., 2021; Deng et al., 2022), similar to our physics-based optimization. However, these methods require significant priors and are limited to single objects, making them less suitable for complex scenes. Our approach also uses MatNet for single-view material property inference but incorporates a physics-based renderer.
@@
Material editing generally follows two categories. The first involves inferring material properties through inverse rendering, then modifying them to achieve edits (Li et al., 2018a; Zhu et al., 2022b; Azinovic et al., 2019). These methods provide strong interpretability, and using a physics-based renderer makes results highly reliable, though realism may sometimes be limited. The second set is neural network-based. Following the success of DM (Rombach et al., 2022), many image editing methods have emerged (Mokady et al., 2023; Wang et al., 2024; Brooks et al., 2023; Shi et al., 2024). Alchemist (Sharma et al., 2024) is a representative example, trained on synthetic data to allow editing of albedo, roughness, metallic properties, and transparency. These methods benefit from the strengths of DM and achieve highly realistic results but have limited interpretability due to their neural network-based nature. Our approach falls into the first category with a solid interpretability, meanwhile, it enables realistic results.
@@
Light estimation is a distinct research field (Wang et al., 2022; Verbin et al., 2024; Park et al., 2020; Swedish et al., 2021; Enyo and Nishino, 2024; Yu et al., 2023a; Tang et al., 2022; Dastjerdi et al., 2023). Some methods represent lighting implicitly (Yao et al., 2022; Zhu et al., 2022b; Kocsis et al., 2024), limiting generalizability, while others require multi-view inputs and scene mesh data (Verbin et al., 2024; Yu et al., 2023a; Park et al., 2020). DPI (Lyu et al., 2023) is related to our approach. This method combines differential rendering with DM for high-quality envmap generation but relies on multi-view NeRF methods for mesh reconstruction and lacks material BRDF data, reducing accuracy. In contrast, our approach uses MatNet’s output for more accurate envmap optimization.
@@
Shadows are crucial for photorealistic rendering. Previous work by Li et al. (Li et al., 2022) emphasized shadow accuracy, using an OptiX-based ray tracer (Parker et al., 2010) to compute shadows for neural network input. However, shadows are often neglected in recent research (Zhu et al., 2022b; Yao et al., 2022; Li et al., 2023); methods based on screen-space or image-based ray tracing lack geometric occlusion and thus fail to generate accurate shadows (Yao et al., 2022; Kocsis et al., 2024; Zhu et al., 2022b). Many neural renderers struggle with shadow generation (Choi et al., 2023; Zhu et al., 2022a), as accurate shadows require physically based light transport, it is worth noting that this limitation has been utilized to identify AI-generated images (Sarkar et al., 2024). Recent advancements in diffusion model-based neural rendering, notably the work presented in DiffusionRenderer (Liang et al., 2025), demonstrate the capacity to generate reasonably realistic shadows without explicit 3D geometry. However, our approach, leveraging 3D geometry, consistently achieves more accurate shadow results compared to DiffusionRenderer. Furthermore, the incorporation of 3D geometry and ray tracing enables physically-based simulation within our framework, a capability currently challenging for diffusion model-based neural rendering techniques.
3. Methods
3.1. Material Prediction Network
@@
Due to the significant success of DM (Rombach et al., 2022) in image generation, recent work used it for the task of material property estimation, achieving impressive results (Kocsis et al., 2024). However, this approach comes with high training costs and requires computationally intensive multi-step sampling during prediction. In comparison, we found that the model architecture for depth estimation (Yang et al., 2024) aligns more closely with the requirements of material prediction. We therefore adopt the DPT architecture (Ranftl et al., 2021), which has been highly successful in depth estimation, for our material prediction task. Different from (Zhu et al., 2022a), we utilize a pretrained DINOv2 (Oquab et al., 2024) encoder for feature extraction. Like previous research in depth estimation (Birkl et al., 2023; Yang et al., 2024), we use two DPT decoders (Ranftl et al., 2021) for the regression of depth and material properties, respectively. Given the advancements in depth estimation, we initialize our training with weights from the pretrained depth model (Yang et al., 2024) to expedite the training process. For the DPT depth decoder, we limit training to the last four layers of the refinenet, while the weights of the DINOv2 encoder remain frozen throughout the training. The following loss function for $\mathcal{L}_{\text{MatNet}}$ is used to optimize the model,
(1) $\mathcal{L}_{\text{MatNet}}=\sum_{i\in\{\mathbf{A},\mathbf{R},\mathbf{M},\mathbf{N},\mathbf{D}\}}\mathcal{L}_{i}\,,$
@@
where the albedo loss $\mathcal{L}_{\mathbf{A}}$ combines the LPIPS perceptual loss and $L_{1}$ loss, while both the roughness loss $\mathcal{L}_{\mathbf{R}}$ and metallic loss $\mathcal{L}_{\mathbf{M}}$ are measured using the $L_{1}$ loss between the predictions and ground truth values. Finally, the $\mathcal{L}_{\mathbf{D}}$ is calculated using SiLogLoss (Bhat et al., 2023; Eigen et al., 2014) to introduce logarithmic error and keep its scale constant to improve the robustness of depth prediction. This loss is defined by
@@
(2) $\mathcal{L}_{\mathbf{D}}=\sqrt{\begin{aligned} &\mathbb{E}_{i\in\text{mask}}\left(\left(\log(t_{i})-\log(p_{i})\right)^{2}\right)\\ &-\lambda\left(\mathbb{E}_{i\in\text{mask}}\left(\log(t_{i})-\log(p_{i})\right)\right)^{2}\,,\end{aligned}}$
@@
(3) $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p},\mathbf{N}_{p},\mathbf{D}_{p}=\text{MatNet}(I)\,,$
@@
Our network is trained on the InteriorVerse (Zhu et al., 2022b) dataset, following the official train/val/test split provided by the dataset to ensure fairness. We apply only random vertical flipping and random 512×512 cropping. The model is trained on one A100 (80GB) for 7 days, optimized using AdamW with a learning rate of 1e-5.
@@
Figure 3. Our single-view mesh reconstruction method used on a test scene. Our method eliminates the white border at the edges. However, since our approach is based on the depth map, it primarily reconstructs the front view of the scene.
@@
3.2. Progressive Differentiable Rendering
We assume a pinhole camera model with a 35-degree field of view, $w\times{h}=512\times 512$ image resolution, and $c=3$ color channels. Our method assumes ambient lighting, represented by an environment map (envmap), and does not explicitly model light sources within the scene. Ignoring emitters in the scene, the rendering equation becomes the reflected radiance equation:
@@ 3.2. Progressive Differentiable Rendering
where $L_{o}(\bm{x},\vec{\omega}_{o})$ is the outgoing radiance at the world space position $\bm{x}$ in the direction $\vec{\omega}_{o}$, $f_{r}(\bm{x},\vec{\omega}_{i},\vec{\omega}_{o})$ is the bidirectional reflectance distribution function (BRDF) at $\bm{x}$, representing the ratio of reflected radiance in the direction $\vec{\omega}_{o}$ to irradiance from the direction $\vec{\omega}_{i}$, while $L_{i}(\bm{x},\vec{\omega}_{i})$ is the incident radiance at $\bm{x}$ from the direction $\vec{\omega}_{i}$, $\vec{n}$ is the surface normal at $\bm{x}$, and $\Omega$ is the hemisphere around $\vec{n}$ covering all directions of incidence. We use Simplified DisneyBRDF to model surface scattering. We let $\mathbf{S}(\bm{x})=\bm{x}_{s}$ denote the transformation from world space to screen space coordinates. Details of coordinate transformation and shading model are in the supplementary material.
@@
We optimize the envmap $\mathbf{E}$ using a position-embedded multi-layer perceptron (MLP), similar to the approach used in NeRF (Mildenhall et al., 2021). Although optimizing $\mathbf{E}$ directly without using a neural network is theoretically possible, this often yields suboptimal results. Standard MLPs are generally insensitive to high-frequency information, making pixel-level optimization of the image difficult; CNNs encounter similar limitations. The envmap is obtained using
(5) $\mathbf{E}=\mathrm{MLP}_{\text{env}}(\gamma(\mathbf{e});\theta_{\text{env}})\,,$
@@
In a differentiable rendering of the geometry $\mathbf{D}_{p},\mathbf{N}_{p}$ based on Monte Carlo integration of Eq. 4, we evaluate the BRDF using $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}$ and the incident radiance using
(6) $L_{i}(\bm{x},\vec{\omega}_{i})=V(\bm{x},\vec{\omega}_{i})L_{\mathbf{E}}(\vec{\omega}_{i})\,,$
where $L_{\mathbf{E}}(\vec{\omega}_{i})$ is the radiance from the envmap in the direction $\vec{\omega}_{i}$ and $V(\bm{x},\vec{\omega}_{i})$ is visibility, which is zero if geometry is found in the direction $\vec{\omega}_{i}$ and one otherwise. Our optimization goal is to find the $\mathbf{E}^{\ast}$ that minimizes the difference between the rendering and the ground truth ($I$):
(7) $\mathbf{E}^{\ast}=\operatorname*{\arg\!\min}_{\mathbf{E}}\mathcal{L_{\mathbf{E}}}(L_{o}(\bm{x},\vec{\omega}_{o}),I)$
where envmap loss $\mathcal{L_{\mathbf{E}}}$ combines $L_{1}$ and $L_{2}$ loss. The network weights $\theta_{\text{env}}$ are discarded after optimization as we only need $\mathbf{E}^{\ast}$.
@@
(8) $\left[\mathbf{A},\mathbf{R},\mathbf{M}\right]=\mathrm{MLP}_{\text{mat}}(\gamma(\mathbf{m});\theta_{\text{mat}})$
(9) $\mathbf{m}=\left[\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}\right]\,,$
@@
(10) $\left[\mathbf{A},\mathbf{R},\mathbf{M}\right]=\zeta\,\tanh(\mathbf{z})+\mathbf{m}\,,$
@@
The rendering process is the same as for the envmap, except that now we use $\bm{E}^{\ast}$ and keep it fixed. Our optimization goal is to find $\mathbf{A}^{\ast},\mathbf{R}^{\ast},\mathbf{M}^{\ast}$ that minimize the difference between the rendered result and the ground truth ($I$):
(11) $\mathbf{A}^{\ast},\mathbf{R}^{\ast},\mathbf{M}^{\ast}=\operatorname*{\arg\!\min}_{\mathbf{A,R,M}}\mathcal{L}_{\text{Mat}}(L_{o}(\bm{x},\vec{\omega}_{o}),I)$
@@
(12) $\mathcal{L}_{\text{Mat}}$ $=\mathcal{L}_{\text{re}}+\delta\mathcal{L}_{\text{cons}}\,,$